The speed of sound in a diatomic gas is measured at 20°C and found to be 215m/s. What is the molecular weight of this gas? What is the most probable composition of this gas?
1
Expert's answer
2020-10-01T10:09:54-0400

The speed of sound in a diatomic gas is


"v=\sqrt{\frac{\gamma RT}{\Mu}}."

For a diatomic gas, with "f=5" degrees of freedom, the heat capacity ratio is


"\gamma=1+\frac{2}{f}=1.4."

Therefore, we can express and calculate the molecular weight M:


"\Mu=\frac{\gamma RT}{v^2}=\frac{1.4\cdot8.314\cdot(273+20)}{215^2}=0.074\text{ kg/mol}."

The most probable composition of this gas is two atoms of chlorine.